# Jon Danielsson London School of Economics Andreas Uthemann

### v2

@@ Abstract
The rapid adoption of artificial intelligence (AI) poses new and poorly understood threats to financial stability. We use a game-theoretic model to analyse the stability impact of AI, finding it amplifies existing financial system vulnerabilities — leverage, liquidity stress and opacity — through superior information processing, common data, speed and strategic complementarities. The consequence is crises become faster and more severe, where the likelihood of a crisis is directly affected by how effectively the authorities engage with AI. In response, we propose the authorities develop their own AI systems and expertise, set up direct AI-to-AI communication, implement automated crisis facilities and monitor AI use.
@@ 1 Introduction
The relationship between artificial intelligence (AI) [^1] Russel (2019) and Norvig and Russell (2021) are particularly useful for a general overview of AI. and financial crises is poorly understood, motivating our work. We use a game theoretic framework to identify how AI impacts on financial stability, finding that AI can both stabilise and destabilise the financial system, depending on the severity of the shocks it faces and how prepared the financial authorities are. As current regulatory frameworks are not adequate for preventing and resolving such crises, we propose ways for the authorities to respond effectively.
@@
While there is no universal definition of AI, the notion of it as a rational maximising agent [^2] One of the definitions given by Norvig and Russell (2021). is particularly useful for the analysis of financial stability since it resonates with methodological approaches taken by most macroprudential investigators. That implies standard frameworks for analysing stability are well suited for studying the impact of AI, not the least game theory, provided it considers AI’s speed and the particular strategic approaches it applies to tasks given to it. To that end, we adopt a GG model and to highlight three key mechanisms through which AI transforms market dynamics: superior information processing that makes signals more precise, common information sources that facilitate coordination and speed advantages that create preemption motives.
@@
Our model yields specific predictions about AI’s impact on crisis dynamics. First, AI makes transitions between stable and crisis states more abrupt. Second, AI increases the likelihood of multiple equilibria and coordinated runs due to ties that have driven financial crises for centuries, ever since the first modern crisis in 1763 (Schnabel and Shin, 2004; Danielsson, 2022): Excessive leverage, liquidity preference during stress and system opacity.
@@
When we apply these theoretical predictions on the impact of AI on crisis dynamics to more practical institutional structure of the financial system, we find four key areas of concern. First, the malicious use of AI — whether to find loopholes, manipulate markets, orchestrate terrorist attacks on financial infrastructure, or enable nation-states to engage in unconventional warfare — poses a serious threat to stabiity. Even a small, coordinated misinformation shock can push collective beliefs past the run threshold and trigger a self-fulfilling crisis.
@@
Second, wrong-way risk refers to a situation where the volume of risky activities increases in line with the riskiness of those activities — and AI gives rise to particularly dangerous forms of wrong-way risk because of how it builds trust. Ultimately, this means that AI reliability is the lowest precisely when it is needed for the most important decisions — the AI wrong-way risk. This may happen when margins set by AI are downward biased, both encouraging risky positions and their risk.
@@
Finally, the increased use of AI affects the speed, intensity, and frequency of crises. When faced with a shock, an AI engine — like a human decision-maker — has a range of options it can deploy in response. But they all boil down to two fundamental choices. Run or stay. Stabilise or destabilise. The consequence of getting it wrong is significant losses and even bankruptcy. AI’s superior ability to both quickly analyse a complex evolving situation and make quick, firm decisions implies it will make crises particularly quick and vicious. When it acts as a crisis amplifier, what might have taken days or weeks to unfold can now happen in minutes or hours. Greater price impact combined with shorter reaction times shortens the adjustment window, shifting the crisis threshold into a state space that was previously safe.
The financial authorities, the central banks and regulators tasked with keeping the system stable, will find it challenging to effectively address the risks arising from those four areas of concern. The problem for the authorities is that they are already losing an arms race with the private sector. To remain effective, we propose that authorities focus on four key areas: developing their own AI engines, establishing AI-to-AI links for direct
So far, most authorities appear to prefer a slow, deliberative and conservative approach to AI. Several have now recognised AI’s implications for financial stability, with recent publications from the IMF (International Monetary Fund, 2024), the BIS (Kiarelly et al., 2024; Aldasoro et al., 2024), the FSB (Financial Stability Board, 2024), the ECB (Leitner et al., 2024), and the Bank of England (Bank of England, 2025).
@@ 2 A Game-Theoretic Framework for AI and Financial Stability
To understand how AI affects financial stability, we develop a formal framework that captures both AI’s unique characteristics and its implications for coordination in financial markets. The key to understanding financial crises lies in how financial institutions optimise — they aim to maximise profits given the acceptable risk. When translating that into how they behave operationally, Roy’s (1952) criterion is particularly useful — stated succinctly, maximising profits subject to not going bankrupt. When financial institutions prioritise survival, their behaviour changes rapidly and drastically. They hoard liquidity and choose the most secure and liquid assets, such as central bank reserves. This leads to bank runs, fire sales, credit crunches and all the other undesirable behaviours associated with crises.
@@
Three characteristics of AI engines particularly impact on how AI affects stability. First, AI systems process data more efficiently than humans, likely giving them more precise signals about market fundamentals and reducing uncertainty. Second, AI systems often rely on similar data, methodologies and especially model architectures. Third, AI can execute transactions much faster than human traders.
The interaction of these mechanisms yields specific predictions about how AI deployment affects financial stability. More precise signals reduce the gray area where fundamentals alone don’t determine individual market participants’ actions, making aggregate behaviour more binary: either no one “runs” (e.g., sells assets or withdraws funding) or everyone does, with transitions becoming sharper and more sudden as fundamentals deteriorate. Common information sources facilitate coordination among AI agents through implicit learning rather than explicit communication, making multiple equilibria more likely, particularly in markets with high AI penetration. Speed advantages create strong preemption motives, with the presence of slower human agents increasing the likelihood of AI-initiated runs. Strategic complementarities amplify these effects through contagion to human agents.
To systematically analyse these dynamics, we adapt the global games framework of Morris and Shin (1998), as it is particularly suitable for analysing coordination problems with strategic uncertainties, where agents make decisions without knowing others’ actions. This framework allows us to develop theoretical predictions about how a putative AI crisis might unfold.
@@ 2.1 Setup
We follow the model setup of Sakovics and Steiner (2012) and apply it to a financial market with a unit mass of traders (AI and human) who must decide whether to “run” or “stay” when faced with a shock that might lead to a crisis. A fraction $\mu$ are AI agents while the rest are human.
The fundamental state $\theta\in[\underline{\theta},\bar{\theta}]$ represents the underlying strength of the financial system, with higher values indicating greater resilience. Let $a\in[0,1]$ denote the mass of investors who decide to run, where, in particular, a crisis occurs if the mass of investors running exceeds the system’s resilience: $a\geq\theta$.
A player’s payoff from choosing to stay given $a$ and $\theta$ is:
@@
A player faces two types of uncertainty when deciding on the optimal strategy: fundamental uncertainty about how vulnerable the system is to runs ($\theta$), and strategic uncertainty about how other market participants will act ($a$).
@@ 2.2 Information Structure and AI Mechanisms
Each agent receives a private signal $x_{i}=\theta+\sigma\eta_{i}$ about the fundamental state $\theta$ and a public signal $y=\theta+\sigma_{p}u$, where $\eta_{i}$ and $u$ are independent random variables with continuous densities and support $[-1/2,1/2]$. A lower $\sigma>0$ implies a more precise private signal, a lower $\sigma_{p}>0$ or more precise public signal. The private signals are independent across players and players only observe their own signal. Agents’ prior about $\theta$ is the uniform distribution on $[\underline{\theta},\bar{\theta}]$ with $\underline{\theta}<\sigma/2$ and $\bar{\theta}>\sigma/2$.
@@ 2.2.1 Mechanism 1: Superior Information Processing
Suppose, for the moment, that there are no public signals. In equilibrium, players follow threshold strategies, running if and only if their private signal falls below a threshold $x^{*}$:
@@
When all players play according to these threshold strategies, a run occurs whenever $\theta<\theta^{\star}$, where $\theta^{\star}$ is the unique value satisfying [^3] At equilibrium, the player with signal $x^{*}=\theta^{*}$ is indifferent between “stay” and “run” because each player best responds to a uniform belief about their opponents’ action choices(Morris and Shin, 2003, see).
@@
AI systems typically process information more efficiently than humans, [^4] Recent evidence suggests that AI systems can achieve superior signal precision in predicting financial crises (Fouliard et al., 2021; Hellwig, 2021; Batsuuri et al., 2024), though this advantage may not extend to unprecedented crisis events with unique characteristics. resulting in more precise private signals (lower $\sigma$). While this does not affect the probability of a successful run — $\theta^{\star}$ does not depend on $\sigma$ — it affects equilibrium behaviour by reducing uncertainty about the fundamental state, making the critical threshold $\theta^{\star}$ for a crisis more sharply defined in terms of the aggregate action. As the precision of the private signal increases (as $\sigma\to 0$), all players coordinate on stay when $\theta\geq\theta^{\star}$ and on run when $\theta<\theta^{\star}$. This leads to fewer “unnecessary” individual run decisions when $\theta>\theta^{\star}$ but increases the completeness of runs when $\theta<\theta^{\star}$.
@@ 2.2.2 Mechanism 2: Common Information Sources
When players observe both private and public signals, they form beliefs about the fundamental by combining these information sources. Hellwig (2002) shows that when public information is sufficiently precise relative to private information, multiple equilibria can emerge. This happens because the public signal $y$ serves as a coordination device, allowing players to synchronise their actions.
AI systems are particularly susceptible to this effect because they draw on common training data and methodologies, share similar model architectures and objective functions and process similar market data in comparable ways. These factors increase the public component of information among AI systems, effectively increasing the weight placed on public signals relative to private signals. Formally, this can be represented as AI agents having a higher relative precision of public to private information ($\sigma/\sigma_{p}$).
The tendency toward common information sources in AI systems creates potential for coordinated behaviour, similar to the portfolio insurance mechanisms analysed by Gennotte and Leland (1990). This coordination can occur without explicit collusion through shared data sources and similar training approaches. (Calvano et al., 2020) provide an example of a market game where Q-learning algorithms engage in collusive pricing using commonly observed price data.
@@ 2.2.3 Mechanism 3: Speed Advantage and Preemption
While humans succeed with probability $p<1$ when choosing to run, AI agents succeed with probability 1. An unsuccessful run yields the same payoff as stay.
@@ 3.1 Malicious Use
Technology has always both helped improve the financial system and provided new ways to destabilise it. For instance, the telegram in 1858 was quickly employed to transmit stock prices across the Atlantic, while algorithmic trading has been implicated in flash crashes (Kirilenko et al., 2017). AI represents a continuation of this pattern, but with important qualitative differences that relate to the three mechanisms described in our framework. The model in the previous section suggests even a small, coordinated misinformation shock can push collective beliefs past the run threshold and trigger a self-fulfilling crisis similar to the political manipulation strategies analyzed in Edmond (2013).
@@
Malicious use of AI in financial markets can occur through several channels. First, bad faith actors can exploit AI’s superior information processing to identify regulatory loopholes or engage in market manipulation. Terrorist groups could leverage AI’s coordination capabilities and speed to orchestrate synchronised attacks on financial infrastructure. Of particular concern is nation-states recruiting AI for unconventional warfare to target vulnerabilities in adversaries’ financial systems while maintaining plausible deniability.
@@
These threats pose a particular challenge to those tasked with defending the financial system — what we term the “defender’s dilemma”. The attackers need only to find a single vulnerability, whereas the defenders must monitor the entire system against potential attacks. The attackers consequently require considerably fewer resources than the defenders; an asymmetry that gets worse with the increasing ability of AI systems.
@@
As a rational maximising agent, AI can also act maliciously without explicit human direction. Scheurer et al. (2024) provide a telling example: a Large Language Model (LLM) was instructed to comply with securities laws and maximise profits. The aim was to investigate how AI would cope with those two potentially conflicting objectives. When given private information, the LLM engaged in illegal insider trading and then lied to its human overseers. This demonstrates a form of model risk that emerges from providing AI with agency. It might find unexpected and potentially harmful paths to attain the goals it is given.
@@
AI’s speed advantage makes it difficult to guard against such behaviour. A potential solution is to maintain humans in-the-loop to make decisions based on AI recommendations, or humans on-the-loop to supervise AI. However, human experts cannot effectively oversee AI operating at machine speeds, and in competitive markets, those who take humans out-of-the-loop can gain the upper hand.
@@
The existing regulatory structure struggles with AI-specific principal-agent problems. Conventional incentive mechanisms that work well with human agents, such as rewards and punishments, lose traction when the “agent” is an optimisation algorithm that cares only about its objective function (Amodei et al., 2016; Leike et al., 2018). Mis-specified goals invite reward hacking and other forms of specification gaming (Krakovna et al., 2020), so the classic one-sided principal-agent problem becomes a two-sided principal-agent-AI problem in which supervisors must monitor both the human intermediaries and the opaque learning system they control.

### v3

@@ Abstract
The rapid adoption of artificial intelligence (AI) poses new and poorly understood threats to financial stability. We use a game-theoretic model to analyse the stability impact of AI, finding that it amplifies existing financial system vulnerabilities — leverage, liquidity stress and opacity — through superior information processing, common data, speed and strategic complementarities. The consequence is crises become faster and more severe, where the likelihood of a crisis is directly affected by how effectively the authorities engage with AI. In response, we propose that the financial authorities develop their own AI systems and expertise, establish direct AI-to-AI communication, implement automated crisis facilities and monitor AI use.
@@ 1 Introduction
The relationship between artificial intelligence (AI) [^1] Russel (2019) and Norvig and Russell (2021) are particularly useful for a general overview of AI. and financial crises is poorly understood, motivating our work. We use a game-theoretic framework to identify how AI impacts financial stability, finding that it can both stabilise and destabilise the financial system, depending on the severity of the shocks it faces and how prepared the financial authorities are. As current regulatory frameworks cannot effectively prevent and resolve such AI-induced crises, we outline how the financial authorities can adapt to AI-driven risks through upgraded supervisory tools, proactive intervention mechanisms and institutional readiness.
@@
While there is no universal definition of AI, the notion of it as a rational maximising agent [^2] One of the definitions given by Norvig and Russell (2021). is particularly useful for the analysis of financial stability since it resonates with methodological approaches taken by most macroprudential investigators. That implies standard frameworks for analysing financial stability are well-suited for studying the impact of AI provided they consider AI’s speed and the particular strategic approaches it applies to tasks given to it. To that end, we adopt a global games model (Morris and Shin, 1998) and highlight three key mechanisms through which AI transforms market dynamics: superior information processing that makes signals more precise, common information sources that facilitate coordination and speed advantages that give rise to preemption motives.
@@
Our model yields specific predictions about AI’s impact on crisis dynamics. First, AI makes transitions between stable and crisis states more abrupt. Second, AI increases the likelihood of multiple equilibria and coordinated runs due to the fundamental factors that have driven financial crises for centuries, ever since the first modern one in 1763 (Schnabel and Shin, 2004; Danielsson, 2022): Excessive leverage, liquidity preference during stress and system opacity.
@@
When we apply these theoretical predictions to a more practical institutional structure of the financial system, we find four key areas of concern. First, the malicious use of AI — whether to find loopholes, manipulate markets, orchestrate terrorist attacks on financial infrastructure, or enable nation-states to engage in unconventional warfare — poses serious threats to stability. Even a small, coordinated misinformation shock can push collective beliefs beyond the tipping point and trigger a self-fulfilling crisis.
@@
Second, wrong-way risk, which refers to a situation in which the volume of risky activities increases in line with the riskiness of those activities. AI gives rise to particularly dangerous forms of wrong-way risk because of how it gains trust, implying that AI reliability is the lowest precisely when it is needed for the most important decisions — the AI wrong-way risk. This may occur when margins set by AI are downward biased, in turn encouraging risky positions and their associated risks.
@@
Finally, the increased use of AI affects the speed, intensity, and frequency of crises. When faced with a shock, an AI engine — like a human decision-maker — has a range of options it can deploy in response. But they all boil down to two fundamental choices. Run or stay. Stabilise or destabilise. The consequences of getting it wrong can be significant, including substantial losses and even bankruptcy. AI’s superior ability to both quickly analyse a complex evolving situation and make quick, firm decisions implies it will make crises particularly quick and vicious. When it acts as a crisis amplifier, what might have taken days or weeks to unfold can now happen in minutes or hours. A greater price impact, combined with shorter reaction times, shortens the adjustment window, shifting the crisis threshold into a state space that was previously considered safe.
The financial authorities — central banks and regulators tasked with maintaining system stability — find it challenging to effectively address the risks arising from these four areas of concern. The problem for the authorities is that they are already losing an arms race with the private sector.
To remain effective, we propose that the authorities focus on several key areas: developing their own AI engines and expertise, establishing AI-to-AI links, implementing automatic facilities, public-private partnerships and particular monitoring of private sector AI use. While most authorities appear to prefer a slow, deliberative and conservative approach to AI, several have now recognised AI’s implications for financial stability; see recent publications from the IMF (International Monetary Fund, 2024), the BIS (Kiarelly et al., 2024; Aldasoro et al., 2024), the FSB (Financial Stability Board, 2024), the ECB (Leitner et al., 2024), and the Bank of England (Bank of England, 2025).
@@ 2 A Game-Theoretic Framework for AI and Financial Stability
To understand how AI affects financial stability, we develop a formal framework that captures both the unique characteristics of AI and its implications for coordination in financial markets. The key to understanding financial crises lies in how financial institutions optimise — they aim to maximise profits given acceptable risk. When translated into operational behaviour, Roy’s (1952) criterion is particularly useful — stated succinctly, maximising profits subject to not going bankrupt. When financial institutions prioritise survival, their behaviour changes rapidly and drastically. They hoard liquidity and choose the most secure and liquid assets, such as central bank reserves. This leads to bank runs, fire sales, credit crunches and all the other undesirable behaviours associated with crises.
@@
Three characteristics of AI engines particularly impact how AI affects stability. First, AI systems process data more efficiently than humans, likely giving them more precise signals about market fundamentals and reducing uncertainty. Second, AI systems often rely on similar data, methodologies and especially model architectures. Third, AI can execute transactions much faster than human traders.
The interaction of these mechanisms yields specific predictions about how the deployment of AI affects financial stability. More precise signals reduce the grey area where fundamentals alone do not determine individual market participants’ actions, making aggregate behaviour more binary: either no one “runs” (e.g., sells assets or withdraws funding) or everyone does, with transitions becoming sharper and more sudden as fundamentals deteriorate. Common information sources facilitate coordination among AI agents through implicit learning rather than explicit communication, making multiple equilibria more likely, particularly in markets with high AI penetration. Speed advantages create strong preemption motives, with the presence of slower human agents increasing the likelihood of AI-initiated runs. Strategic complementarities amplify these effects through contagion to human agents.
To systematically analyse these dynamics, we adapt the global game’s framework of Morris and Shin (1998), as it is particularly suitable for analysing coordination problems with strategic uncertainties, where agents make decisions without knowing others’ actions. This framework allows us to develop theoretical predictions about how a putative AI crisis might unfold.
@@ 2.1 Setup
We adapt Sakovics and Steiner’s (2012) model setup to a financial market with a unit mass of investors (AI and human) who must decide whether to “run” or “stay” when faced with a shock that might lead to a crisis.
There is a continuum of investors of unit mass indexed by $i\in[0,1]$. A fraction $\mu$ of investors are AI while the rest are human. The fundamental state $\theta\in[\underline{\theta},\bar{\theta}]$ represents the underlying strength of the financial system, with higher values indicating greater resilience. Let $a\in[0,1]$ denote the mass of investors who decide to run, where, in particular, a crisis occurs if the mass of investors running exceeds the system’s resilience: $a\geq\theta$.
An investor’s payoff from choosing to stay given $a$ and $\theta$ is:
@@
An investor faces two types of uncertainty when deciding on the optimal strategy: fundamental uncertainty about how vulnerable the system is to runs ($\theta$) and strategic uncertainty about how other market participants will act ($a$).
@@ 2.2 Information Structure and AI Mechanisms
Each investor receives two noisy signals about the fundamental state $\theta$: a private signal $x_{i}=\theta+\sigma\eta_{i}$ and a public signal $y=\theta+\sigma_{p}u$. All investors observe $y$, but only investor $i$ observes $x_{i}$. The common noise term $u$ and the private noise terms $\eta_{i}$ for all investors $i\in[0,1]$ are random variables with continuous densities and support $[-1/2,1/2]$ and are mutually independent. $1/\sigma^{2}$ is the precision of the private signal and $1/\sigma_{p}^{2}$ is the precision of the public signal. A higher precision implies a more accurate signal about the fundamental. Investors’ prior about $\theta$ is the uniform distribution on $[\underline{\theta},\bar{\theta}]$ with $\underline{\theta}<\sigma/2$ and $\bar{\theta}>\sigma/2$.
@@ 2.2.1 Mechanism 1: Superior Information Processing
Suppose for the moment that there are no public signals. In equilibrium, investors follow threshold strategies, running if and only if their private signal falls below a threshold $x^{*}$:
@@
When all investors act according to these threshold strategies, a run occurs whenever $\theta<\theta^{\star}$, where $\theta^{\star}$ is the unique value satisfying [^3] At equilibrium, the investor with signal $x^{*}=\theta^{*}$ is indifferent between “stay” and “run” because each investor best-responds to a uniform belief about their opponents’ action choices (Morris and Shin, 2003, see).
@@
AI systems typically process information more efficiently than humans, [^4] Recent evidence suggests that AI systems can achieve superior signal precision in predicting financial crises (Fouliard et al., 2021; Hellwig, 2021; Batsuuri et al., 2024), though this advantage may not extend to unprecedented crisis events with unique characteristics. resulting in more precise private signals (lower $\sigma$). While this does not affect the probability of a successful run — $\theta^{\star}$ does not depend on $\sigma$ — it affects equilibrium behaviour by reducing uncertainty about the fundamental state, making the critical threshold $\theta^{\star}$ for a crisis more sharply defined in terms of the aggregate action. As the precision of the private signal increases (as $\sigma\to 0$), all investors coordinate on stay when $\theta\geq\theta^{\star}$ and on run when $\theta<\theta^{\star}$. This leads to fewer “unnecessary” individual run decisions when $\theta>\theta^{\star}$ but increases the completeness of runs when $\theta<\theta^{\star}$.
@@ 2.2.2 Mechanism 2: Common Information Sources
When investors observe both private and public signals, they form beliefs about the fundamentals by combining these information sources. Hellwig (2002) shows that when public information is sufficiently precise relative to private information, multiple equilibria can emerge. This happens because the public signal $y$ serves as a coordination device, allowing investors to synchronise their actions.
AI systems are particularly susceptible to this synchronisation effect because they draw on common training data and methodologies, share similar model architectures and objective functions and process similar market data in comparable ways. These factors enhance the public component of information in AI systems, thereby increasing the weight placed on public signals relative to private signals. Formally, this can be represented as AI agents having a higher relative precision of public to private information ($\sigma/\sigma_{p}$).
The tendency toward common information sources in AI systems creates a potential for coordinated behaviour, similar to the portfolio insurance mechanisms analysed by Gennotte and Leland (1990). This coordination can occur without explicit collusion through shared data sources and similar training approaches. Calvano et al. (2020) provide an example of a market game where Q-learning algorithms engage in collusive pricing using commonly observed price data.
@@ 2.2.3 Mechanism 3: Speed Advantage and Preemption
While humans succeed with probability $p<1$ when choosing to run, AI agents succeed with probability 1. An unsuccessful run yields the same payoff as a stay.
@@ 3.1 Malicious Use
Technology has always both helped improve the financial system and provided new ways to destabilise it. For instance, the telegram in 1858 was quickly employed to transmit stock prices across the Atlantic, while algorithmic trading has been implicated in flash crashes (Kirilenko et al., 2017). AI is the same but with important qualitative differences that relate to the three mechanisms described in our framework. The model in the previous section suggests that even a small, coordinated misinformation shock can push collective beliefs past the run threshold and trigger a self-fulfilling crisis, similar to the political manipulation strategies analysed in Edmond (2013).
@@
Malicious use of AI in financial markets can occur through several channels. First, bad-faith actors can exploit AI’s superior information processing capabilities to identify regulatory loopholes or engage in market manipulation. Terrorist groups could leverage AI’s coordination capabilities and speed to orchestrate synchronised attacks on financial infrastructure. Of particular concern is nation-states recruiting AI for unconventional warfare to target vulnerabilities in adversaries’ financial systems while maintaining plausible deniability.
@@
These threats pose a particular challenge to those tasked with defending the financial system — what we term the “defender’s dilemma”. The attackers need only to find a single vulnerability, whereas the defenders must monitor the entire system against potential attacks. The attackers consequently require considerably fewer resources than the defenders, an asymmetry that gets worse with the increasing ability of AI systems.
@@
As a rational maximising agent, AI can also act maliciously without explicit human direction. Scheurer et al. (2024) provide a telling example: a large language model (LLM) was instructed to comply with securities laws and maximise profits. The aim was to investigate how AI would address these two potentially conflicting objectives. When given private information, the LLM engaged in illegal insider trading and then lied to its human overseers. This demonstrates a form of model risk that emerges from providing AI with agency. It might find unexpected and potentially harmful paths to attain the goals it is given.
@@
AI’s speed advantage makes it difficult to guard against such behaviour. A potential solution is to maintain humans in-the-loop to make decisions based on AI recommendations, or humans on-the-loop to supervise AI. However, human experts cannot effectively oversee AI operating at machine speeds, and in competitive markets, those who remove humans from the loop can gain the upper hand.
@@
The existing regulatory structure struggles with AI-specific principal-agent problems. Conventional incentive mechanisms which work well with human agents, such as rewards and punishments, lose traction when the agent is an optimisation algorithm that cares only about its objective function (Amodei et al., 2016; Leike et al., 2017). Mis-specified goals invite reward hacking and other forms of specification gaming (Krakovna et al., 2020), so the classic one-sided principal-agent problem becomes a two-sided principal-agent-AI problem in which supervisors must monitor both the human intermediaries and the opaque learning system they control.